# The linearization approach to the Calderón problem revisited: reconstruction via the Born approximation

### v1

@@ 2.1. Linearized Calderón problem
    and $P_{0}$ is nothing but the Poisson integral, which maps continuously $L^{2}(\partial\Omega)$ in $H^{1/2}(\Omega)$ (see, for instance, [37, Section 7.12]), it follows that
@@ 2.2. The Born approximation in the Calderón problem
To prove the converse, assume that a $\gamma^{\mathrm{B}}_{\sigma}\in W^{2,\infty}(\Omega)+\mathcal{E}^{\prime}(\Omega)$ exists such that $\mathrm{d}\Phi_{\sigma}(\gamma^{\mathrm{B}}_{\sigma})=\Lambda_{\gamma}$. We first show that $\gamma^{\mathrm{B}}_{\sigma}|_{\partial\Omega}=\gamma|_{\partial\Omega}$; to do this, we use a construction due to Brown [8] (see also [21, Theorem A1]). Fix $x\in\partial\Omega$ and consider the bounded sequence $(f_{N}^{x})_{N\in\mathbb{N}}$ in $H^{1/2}(\partial\Omega)$ defined in [21, Theorem A1]. This result, combined with (2.1) and (2.2), implies that:
@@ 2.4. Existence of the Born approximation in the radial case
$\left\langle\overline{f},\mathrm{d}\Phi_{1}(\gamma)g\right\rangle_{H^{1/2}\times H^{-1/2}}=\ell(2\ell+d-2)\frac{\left\langle f,g\right\rangle_{L^{2}(\mathbb{S}^{d-1})}}{|\mathbb{S}^{d-1}|}\int_{\mathbb{B}^{d}}\gamma(x)|x|^{2\ell-2}\,\mathrm{d}x.$
@@
In [17, Theorem 1] it is proved that there exist a real and radial function $\gamma^{\mathrm{B}}_{1}\in W^{2,1}(\mathbb{B}^{d})$ such that the identity above holds. This proves that (1.8) holds for $\kappa=0$. In fact, is the unique solution in $L^{\infty}(\mathbb{B}^{d})+\mathcal{E}^{\prime}(\mathbb{B}^{d})$ by Theorem 2.7. The identity of the traces at the boundary and that $\gamma^{\mathrm{B}}_{1}\in\mathcal{C}^{1}(\overline{\mathbb{B}^{d}}\setminus\{0\})$ also follow from [17, Theorem 1]. This proves the theorem for $\kappa=0$. ∎
@@
In [32, Theorems 1 and 3.1], it is shown that $\mathrm{d}\Phi^{\rm S}_{-\kappa}$ extends to the space $\mathcal{B}_{d}:=\mathcal{E}^{\prime}_{\mathrm{rad}}(\mathbb{B}^{d})+L^{1}_{\mathrm{rad}}(\mathbb{B}^{d})$ and that for every $q\in L^{\infty}(\mathbb{B}^{d})$ one can find a radial distribution $q^{\rm B}_{-\kappa}\in\mathcal{B}_{d}$ such that $\mathrm{d}\Phi^{\rm S}_{-\kappa}(q^{\rm B}_{-\kappa}+\kappa)=\Lambda^{\rm S}_{q}-\Lambda^{\rm S}_{-\kappa}$. This is stated, specifically in [32, Equation (1.15)], which in our notation [^6] The potential in [32] corresponds in our notation to $q+\kappa$, and the DtN map $\Lambda_{q,\kappa}$ in [32] corresponds here to $\Lambda_{q+\kappa}^{\mathrm{S}}$. reads
@@
The fact that the distribution $q^{\rm B}_{-\kappa}$ is a function outside the origin and that $q-q^{\rm B}_{-\kappa}\in\mathcal{C}(\mathbb{B}^{d}\setminus\{0\})$ it is proved in [32, Theorem 2].
It remains to prove that $q=q(\gamma)$ implies that $q^{\rm B}_{-\kappa}\in L^{1}(\mathbb{B}^{d})$. For $\kappa=0$, this is proved in [17, Proposition 2.1]. The case $\kappa\neq 0$ is essentially contained in the works [32, 17], though it is not stated explicitly in them. Here we outline the chain of arguments that leads to proving that $q^{\rm B}_{-\kappa}\in L^{1}(\mathbb{B}^{d})$:
  • Let $Q\in L^{1}_{\mathrm{loc}}(\mathbb{R}_{+})$ such that $Q(-\log|x|)=q(x)|x|^{2}$. Since $q$ arises from a conductivity [17, Theorem 3.2 and Lemma 3.9] imply that [17, Identity (3.13)] holds for all $z\geq\nu_{d}$.
  • The identity [17, Identity (3.13)] is the same as [32, Equation (4.11)], so, again, it holds for all $z\geq z_{Q}:=\nu_{d}$ and $d\geq 2$ (in particular, in $d=2$ this implies that the assumption $z_{Q}>0$ in [32, Equation (4.11)] can be relaxed to $z_{Q}=0$.)
  • On the other hand, the conclusion of [32, Lemma 4.3] holds with $z_{\kappa}=0$ since $\kappa<\lambda_{0,1}^{2}$.
  • As a consequence of the two previous steps, the conclusion of [32, Proposition 4.6] is valid with $\ell_{q}:=\max(z_{Q},z_{\kappa})-\nu_{d}=0$ (the key here is that one can take $z_{Q}=\nu_{d}$ in [32, Proposition 4.6] instead of the value of $z_{Q}$ given by [32, Identity (4.12)].). In particular, [32, Identity (4.19)] holds for a function $q^{s}_{\kappa}\in L^{1}(\mathbb{B}^{d})$. This identity is equivalent to (2.27) taking $q^{\rm B}_{-\kappa}=q^{s}_{\kappa}-\kappa$, which proves that $q^{\rm B}_{-\kappa}\in L^{1}(\mathbb{B}^{d})$.
@@ 3.1. The Born approximation and the complex moment problem when $\kappa=0$
The following lemma allows us to interpret the matrix elements of $\mathrm{d}\Phi_{1}(\gamma)$ and $\mathrm{d}\Phi_{-\kappa}^{\rm S}(q)$ as complex moments of $\gamma$ and $q$, respectively.
@@
$\left\langle\overline{e_{\ell}},\mathrm{d}\Phi_{1}(\gamma)e_{m}\right\rangle_{H^{1/2}\times H^{-1/2}}=2\ell m\mu_{\ell-1,m-1}^{0}[\gamma].$
Also
$\left\langle\overline{e_{\ell}},\mathrm{d}\Phi_{0}^{\rm S}(q)e_{m}\right\rangle_{H^{1/2}\times H^{-1/2}}=\mu_{\ell,m}^{0}[q],\qquad\ell,m\in\mathbb{N}.$
@@ 3.2. Connection with moment problems in the general case
(see [32, Lemma 2.1] for a proof). This finishes the proof of the lemma. ∎
@@ 3.3. Uniqueness of the complex moment problems
From the generating function for the Bessel coefficients [39, Section 2.1], and the identity $I_{n}(z)=(-i)^{n}J_{n}(iz)$ for $n\in\mathbb{Z}$, we obtain
@@
The result follows from (see [39, Section 8.1])
@@
4. Numerical results
@@ 4.1. Direct problem method
We aim to compute the matrix elements $\left\langle\overline{e_{\ell}},\Lambda_{\gamma}e_{m}\right\rangle_{H^{1/2}\times H^{-1/2}}$ from a given $\gamma$. In order to do this, we first need to solve a discretized version of (1.1) with $f=e_{m}$. We embed in $\mathbb{D}$ a spiderweb graph that has the positive, second kind Chebyshev nodes of order $2N_{r}-1$ in the radial direction and an even number $N_{\theta}$ of Fourier nodes in the angular direction. Furthermore, we assign to each direction its corresponding spectral differentiation matrix. Following [38, Chapter 11], we use these nodes and differentiation matrices to transform (1.1) into a linear matrix equation, which we then solve. The projection over $e_{\ell}$ is then done with a trapezoidal integration rule. Since we use $N_{\theta}$ Fourier nodes, it only makes sense to compute the matrix elements for which $\left\lvert l\right\rvert,\left\lvert m\right\rvert<N_{\theta}/2$.
@@ 4.2. Inverse problem method
If the L-curve has a sharp corner, then we choose the $\lambda$ at which the curvature of the L-curve is maximal (we actually maximize a simplified version of the curvature introduced in [28, Lemma 1.1]), otherwise we choose the $\lambda$ at which the curvature vanishes.

### v2

@@ 2.1. Linearized Calderón problem
    and $P_{0}$ is nothing but the Poisson integral, which maps continuously $L^{2}(\partial\Omega)$ in $H^{1/2}(\Omega)$ (see, for instance, [44, Section 7.12]), it follows that
@@ 2.2. The Born approximation in the Calderón problem
To prove the converse, assume that a $\gamma^{\mathrm{B}}_{\sigma}\in W^{2,\infty}(\Omega)+\mathcal{E}^{\prime}(\Omega)$ exists such that $\mathrm{d}\Phi_{\sigma}(\gamma^{\mathrm{B}}_{\sigma})=\Lambda_{\gamma}$. We first show that $\gamma^{\mathrm{B}}_{\sigma}|_{\partial\Omega}=\gamma|_{\partial\Omega}$; to do this, we use a construction due to Brown [8] (see also [22, Theorem A1]). Fix $x\in\partial\Omega$ and consider the bounded sequence $(f_{N}^{x})_{N\in\mathbb{N}}$ in $H^{1/2}(\partial\Omega)$ defined in [22, Theorem A1]. This result, combined with (2.1) and (2.2), implies that:
@@ 2.4. Existence of the Born approximation in the radial case
(2.27) $\left\langle\overline{f},\mathrm{d}\Phi_{1}(\gamma)g\right\rangle_{H^{1/2}\times H^{-1/2}}=\ell(2\ell+d-2)\frac{\left\langle f,g\right\rangle_{L^{2}(\mathbb{S}^{d-1})}}{|\mathbb{S}^{d-1}|}\int_{\mathbb{B}^{d}}\gamma(x)|x|^{2\ell-2}\,\mathrm{d}x.$
@@
In [18, Theorem 1] it is proved that there exist a real and radial function $\gamma^{\mathrm{B}}_{1}\in W^{2,1}(\mathbb{B}^{d})$ such that the identity above holds. This proves that (1.8) holds for $\kappa=0$. In fact, is the unique solution in $L^{\infty}(\mathbb{B}^{d})+\mathcal{E}^{\prime}(\mathbb{B}^{d})$ by Theorem 2.7. The identity of the traces at the boundary and that $\gamma^{\mathrm{B}}_{1}\in\mathcal{C}^{1}(\overline{\mathbb{B}^{d}}\setminus\{0\})$ also follow from [18, Theorem 1]. This proves the theorem for $\kappa=0$. ∎
@@
In [35, Theorems 1 and 3.1], it is shown that $\mathrm{d}\Phi^{\rm S}_{-\kappa}$ extends to the space $\mathcal{B}_{d}:=\mathcal{E}^{\prime}_{\mathrm{rad}}(\mathbb{B}^{d})+L^{1}_{\mathrm{rad}}(\mathbb{B}^{d})$ and that for every $q\in L^{\infty}(\mathbb{B}^{d})$ one can find a radial distribution $q^{\rm B}_{-\kappa}\in\mathcal{B}_{d}$ such that $\mathrm{d}\Phi^{\rm S}_{-\kappa}(q^{\rm B}_{-\kappa}+\kappa)=\Lambda^{\rm S}_{q}-\Lambda^{\rm S}_{-\kappa}$. This is stated, specifically in [35, Equation (1.15)], which in our notation [^6] The potential in [35] corresponds in our notation to $q+\kappa$, and the DtN map $\Lambda_{q,\kappa}$ in [35] corresponds here to $\Lambda_{q+\kappa}^{\mathrm{S}}$. reads
@@
The fact that the distribution $q^{\rm B}_{-\kappa}$ is a function outside the origin and that $q-q^{\rm B}_{-\kappa}\in\mathcal{C}(\mathbb{B}^{d}\setminus\{0\})$ it is proved in [35, Theorem 2].
It remains to prove that $q=q(\gamma)$ implies that $q^{\rm B}_{-\kappa}\in L^{1}(\mathbb{B}^{d})$. For $\kappa=0$, this is proved in [18, Proposition 2.1]. The case $\kappa\neq 0$ is essentially contained in the works [35, 18], though it is not stated explicitly in them. Here we outline the chain of arguments that leads to proving that $q^{\rm B}_{-\kappa}\in L^{1}(\mathbb{B}^{d})$:
  • Let $Q\in L^{1}_{\mathrm{loc}}(\mathbb{R}_{+})$ such that $Q(-\log|x|)=q(x)|x|^{2}$. Since $q$ arises from a conductivity [18, Theorem 3.2 and Lemma 3.9] imply that [18, Identity (3.13)] holds for all $z\geq\nu_{d}$.
  • The identity [18, Identity (3.13)] is the same as [35, Equation (4.11)], so, again, it holds for all $z\geq z_{Q}:=\nu_{d}$ and $d\geq 2$ (in particular, in $d=2$ this implies that the assumption $z_{Q}>0$ in [35, Equation (4.11)] can be relaxed to $z_{Q}=0$.)
  • On the other hand, the conclusion of [35, Lemma 4.3] holds with $z_{\kappa}=0$ since $\kappa<\lambda_{0,1}^{2}$.
  • As a consequence of the two previous steps, the conclusion of [35, Proposition 4.6] is valid with $\ell_{q}:=\max(z_{Q},z_{\kappa})-\nu_{d}=0$ (the key here is that one can take $z_{Q}=\nu_{d}$ in [35, Proposition 4.6] instead of the value of $z_{Q}$ given by [35, Identity (4.12)].). In particular, [35, Identity (4.19)] holds for a function $q^{s}_{\kappa}\in L^{1}(\mathbb{B}^{d})$. This identity is equivalent to (2.28) taking $q^{\rm B}_{-\kappa}=q^{s}_{\kappa}-\kappa$, which proves that $q^{\rm B}_{-\kappa}\in L^{1}(\mathbb{B}^{d})$.
@@ 3.1. The Born approximation and the complex moment problem when $\kappa=0$
The following lemma allows us to interpret the matrix elements of $\mathrm{d}\Phi_{1}(\gamma)$ and $\mathrm{d}\Phi_{0}^{\rm S}(q)$ as complex moments of $\gamma$ and $q$, respectively.
@@
$\left\langle\overline{e_{\ell}},\mathrm{d}\Phi_{1}(\gamma)e_{m}\right\rangle_{H^{1/2}\times H^{-1/2}}$ $=2\ell m\mu_{\ell-1,m-1}^{0}[\gamma],$
$\left\langle\overline{e_{\ell}},\mathrm{d}\Phi_{0}^{\rm S}(q)e_{m}\right\rangle_{H^{1/2}\times H^{-1/2}}$ $=\mu_{\ell,m}^{0}[q].$
@@ 3.2. Connection with moment problems in the general case
(see [35, Lemma 2.1] for a proof). This finishes the proof of the lemma. ∎
@@ 3.3. Uniqueness of the complex moment problems
From the generating function for the Bessel coefficients [46, Section 2.1], and the identity $I_{n}(z)=(-i)^{n}J_{n}(iz)$ for $n\in\mathbb{Z}$, we obtain
@@
The result follows from (see [46, Section 8.1])
@@
4. Numerical analysis of the Born approximation
@@ 4.1. Direct problem method
We aim to compute the matrix elements $\left\langle\overline{e_{\ell}},\Lambda_{\gamma}e_{m}\right\rangle_{H^{1/2}\times H^{-1/2}}$ from a given $\gamma$. In order to do this, we first need to solve a discretized version of (1.1) with $f=e_{m}$. We embed in $\mathbb{D}$ a spiderweb graph that has the positive, second kind Chebyshev nodes of order $2N_{r}-1$ in the radial direction and an even number $N_{\theta}$ of Fourier nodes in the angular direction. Furthermore, we assign to each direction its corresponding spectral differentiation matrix. Following [45, Chapter 11], we use these nodes and differentiation matrices to transform (1.1) into a linear matrix equation, which we then solve. The projection over $e_{\ell}$ is then done with a trapezoidal integration rule. Since we use $N_{\theta}$ Fourier nodes, it only makes sense to compute the matrix elements for which $\left\lvert l\right\rvert,\left\lvert m\right\rvert<N_{\theta}/2$.
@@ 4.2. Inverse problem method
If the L-curve has a sharp corner, then we choose the $\lambda$ at which the curvature of the L-curve is maximal (we actually maximize a simplified version of the curvature introduced in [30, Lemma 1.1]), otherwise we choose the $\lambda$ at which the curvature vanishes.